OpenStudy (anonymous):

Please help me with these two last problems. No more after this today! I mean it, please help me out, i have less than an hour left!!

OpenStudy (lilg132):

whats the question

OpenStudy (anonymous):

\[4x=\sqrt{2x+3}\]

OpenStudy (lilg132):

4x^2 = 2x + 3 4x^2 - 2x - 3 = 0 use quadratic formula

OpenStudy (anonymous):

can i use factoring?

OpenStudy (anonymous):

do you want to learn how to do it or do you just want the answer?

OpenStudy (anonymous):

no the wolf thing doesnt work lown, but thanks

OpenStudy (anonymous):

sure it does

OpenStudy (anonymous):

Type this into wolfram alpha: Solve[4x == Sqrt[2x + 3], x]

OpenStudy (anonymous):

typing in what you put up there worked, thanks :)

hero (hero):

4x^2 = 2x + 3 4x^2-2x - 3 = 0 Yup, Quad formula

OpenStudy (lilg132):

lol so back to JUST getting the answers and not learning lol

OpenStudy (anonymous):

that's all she wants, i'd rather all the leeches just go to wolfram so that people here can actually learn

hero (hero):

lilg, you already have something to work on

OpenStudy (anonymous):

It also has the advantage of making fewer people give trivial answers if the people with trivial questions can get them solved by computer anyway

OpenStudy (lilg132):

hero im waiting for a reply from you

OpenStudy (anonymous):

I need to know this eventually ya know, i have to take my own tests etc, BUT I need these now. And i dont have time to settle down and try them out, and i am without paper, and its awfully hard to work something out paper-less Also, don't think that just because i have a hard time with math means im unintelligent or a 'leech' i appreciate all the help i get, it really makes a different. Finally, the only reason i need a dumb math class is because im getting a bachelors of science degree, but i am an artist, and my career really doesnt demand it, only my degree.

OpenStudy (anonymous):

Well somday if you want tutoring, i'm sure people here are happy to give it to you, but if you just want answers then you can get it from the automatic wolfram web site.

OpenStudy (lilg132):

no one called you unintelligent just lazy and theres always notepad or paint to work with if you really wanted to :p anyways getting JUST the answers isnt going to help you so not any of our loss

OpenStudy (anonymous):

Thanks again ktklown It just felt that way lilg132